OpenStudy (anonymous):

8x–2y=4 4x+5y=–18 How do you solve using the substitution method?

OpenStudy (anonymous):

First we'll need to isolate one of the variables. Lets start with the equation 8x - 2y = 4 8x - 2y = 4 Lets subtract 8x from both sides: -2y = 4 - 8x Now we'll divide both sides by -2: y = -2 + 4x Now we have the (y) variable isolated in terms of (x), so lets substitute our equation in for (y) in the second equation: 4x + 5y = -18 We found out that y = -2 + 4x, so lets plug that in for the (y) above: 4x + 5( -2 + 4x ) = -18 4x - 10 + 20x = -18 24x = 18 x = 3/4 Now that we know what (x) equals to, lets plug into our equation y = -2 + 4x to find out (y): y = -2 + 4x y = -2 + 4( 3/4) y = -2 + 3 y = 1 so x =3/4 and y =1